# Imagination Augmented Generation: Learning to Imagine Richer Context for Question Answering over Large Language Models

### v2

@@ A.5.2 OOD Results
A.5.3 Zero-Shot results Best_Subspan EM
@@
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with an increase in document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text imagination during the compression phase.
@@
A.5.4 OOD and Ablation Experiment Results
@@
T5 Roberts et al. (2020a) 220M 0 25.9* 23.8* 27.9*
T5-l Roberts et al. (2020a) 770M 0 28.5* 28.7* 30.6*
T5-xl 3b 0 28.30 33.92 34.43
LoRA-Base Hu et al. (2021) 220M 0 16.17 21.16 26.38
LoRA-l Hu et al. (2021) 770M 0 17.70 23.87 29.13
@@
DPR Karpukhin et al. (2020) 110M 100 41.5* 56.8* 41.1*
RAG Lewis et al. (2020) 400M 10 44.5* 56.1* 45.2*
FiD Izacard and Grave (2021) 220M 100 48.2* 65.0* 46.71
FiD-l 770M 100 51.4* 67.6* 50.52
FiD-xl 3b 20 55.18 72.92 52.85
FiD-l 770M 10 46.7* 61.9* 48.1*
FiD-xl 3b 10 50.1* 66.3* 50.8*
@@
GENREAD-l Yu et al. (2023) 770M 10$\dagger$ 40.3* 67.8* 51.5*
GENREAD-xl 3b 10$\dagger$ 42.6* 69.6* 52.6*
# Our proposed method
LoRA-Base 220M 1$\dagger$ 34.51 54.05 32.28
LoRA-l 770M 1$\dagger$ 40.05 62.81 43.70
LoRA-xl 3b 1$\dagger$ 44.15 66.92 48.23
IMcQA 220M 1$\dagger$ 40.14 60.75 41.73
IMcQA-l 770M 1$\dagger$ 42.32 65.48 45.28
IMcQA-xl 3b 1$\dagger$ 46.51 68.38 50.45
IMcQA 220M 10 47.01 64.95 46.36
IMcQA-l 770M 10 49.92 69.67 51.52
IMcQA-xl 3b 5$\ddagger$ 50.87 70.34 52.78
Table 8: Full QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that explicit documentation was not utilized; The latter three parts utilize explicit augmented documents. The best results are in bold. * means that those results are from existing papers, $\dagger$ denotes that the number of documents is generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation).

### v3

@@ C.2 OOD Results
C.3 Zero-Shot Results
@@
However, the model did exhibit a weak adherence to instructions, often failing to output the exact answer. Remarkably, Llama2-13B displayed a decline in EM with increased document length on the WQ dataset, whereas the Best_Subspan_EM value augmented. Contrarily, our method excelled in extracting key information by using text imagination during the compression phase.
Model NQ TriviaQA WebQ
LLaMA-2-7B
Zero-shot 8.6 14.5 2.6
DPR + ICL 18.3 32.5 15.6
DPR + RECITE Sun et al. (2023) 16.8 43.9 24.8
DPR + HICL Wang et al. (2024) 25.1 47.5 28.1
DPR + ImcQA (Ours) 33.7 44.5 31.9
Table 8: Zero-shot results of Llama2-7B
C.4 OOD and Ablation Experiment Results
C.5 Error Analysis
Using LLM as a knowledge base inevitably leads to hallucinations, which is a significant area of research in LLM development. In our quality analysis, we sampled 100 generated documents. As shown in Table 10, we found that hallucinations occurred with a probability of 4%, while the occurrence of meaningless text, such as repeated values, was 6%. Consequently, the impact of hallucinations in our method is relatively minor.
@@ C.5 Error Analysis
Hallucinations Meaningless
4% 6%
Question: When is the next Deadpool movie being released? Question: Who got the first Nobel Prize in Physics?
Document: "Deadpool (film) Deadpool is a 2016 American superhero film based on the Marvel Comics character of the same name, produced by Marvel Studios and distributed by Walt Disney Studios Motion Pictures. Document: The Nobel Prize is not a prize in itself.
Correct answer: May 18, 2018 Correct answer: Wilhelm Conrad Röntgen
Table 10: Hallucinations and Meaningless Analysis.
@@
T5 Roberts et al. (2020a) 220M 0 25.9 23.8 27.9
T5-l Roberts et al. (2020a) 770M 0 28.5 28.7 30.6
T5-xl Roberts et al. (2020a) 3b 0 28.30 33.92 34.43
LoRA-Base 220M 0 16.17 21.16 26.38
LoRA-l 770M 0 17.70 23.87 29.13
@@
DPR Karpukhin et al. (2020) 110M 100 41.5 56.8 41.1
RAG Lewis et al. (2020) 400M 10 44.5 56.1 45.2
FiD Izacard and Grave (2021) 220M 100 48.2 65.0 46.71
FiD-l Izacard and Grave (2021) 770M 100 51.4 67.6 50.52
FiD-xl Izacard and Grave (2021) 3b 20 55.18 72.92 52.85
FiD-l Izacard and Grave (2021) 770M 10 46.7 61.9 48.1
FiD-xl Izacard and Grave (2021) 3b 10 50.1 66.3 50.8
EAR-l Chuang et al. (2023) 770M 10 39.6 60.0 -
EAR-xl Chuang et al. (2023) 3b 10 42.3 64.6 -
RFiD-l Wang et al. (2023a) 770M 10 48.3 63.4 -
RFiD-xl Wang et al. (2023a) 3b 10 50.5 67.8 -
FILCO-xl Wang et al. (2023d) 3b 1 44.7 59.0 -
IMcQA (Ours) 220M 10 47.01 64.95 46.36
IMcQA-l (Ours) 770M 10 49.92 69.67 51.52
IMcQA-xl (Ours) 3b 5 50.87 70.34 52.78
IMcQA-l (Ours) 770M 30 53.1 70.5 52.0
@@
GENREAD-l (sampling) Yu et al. (2023) 770M 10 40.3 67.8 51.5
GENREAD-l (clustering) Yu et al. (2023) 770M 10 43.5 70.2 53.5
GENREAD-xl (sampling) Yu et al. (2023) 3b 10 42.6 69.6 52.6
GENREAD-xl (clustering) Yu et al. (2023) 3b 10 45.6 71.6 54.4
IMcQA (Ours) 220M 10 46.22 66.70 51.43
IMcQA-l (Ours) 770M 10 48.83 70.85 54.52
IMcQA-xl (Ours) 3b 5 49.23 72.18 55.39
# Imagination Augmented Generation (Ours)
LoRA-Base 220M 1 34.51 54.05 32.28
LoRA-l 770M 1 40.05 62.81 43.70
LoRA-xl 3b 1 44.15 66.92 48.23
IMcQA 220M 1 40.14 60.75 41.73
IMcQA-l 770M 1 42.32 65.48 45.28
IMcQA-xl 3b 1 46.51 68.38 50.45
Table 12: Full QA performances (%) of different methods on three datasets. The first part (closed-book setting) indicates that explicit documentation was not utilized; The latter three parts utilize explicit augmented documents. The best results are in bold. * means that those results are from existing papers, denotes that the number of documents is generated ($\ddagger$ indicates that the number of documents is reduced due to insufficient memory for distillation).